### v1

# Injecting Domain-Specific Knowledge into Large Language Models: A Comprehensive Survey

Abstract
[…]
1 Introduction
Large Language Models (LLMs) have achieved extraordinary success across various tasks, showcasing remarkable capabilities in reasoning, knowledge representation, and decision-making. However, despite their impressive performance in general-purpose applications, many specialized domains, such as healthcare, chemistry, and legal analysis, demand the integration of domain-specific knowledge to achieve high accuracy and reliability. To address this challenge, researchers have explored methods to enhance LLMs through external or embedded domain expertise, a process often referred to as knowledge injection. This approach aims to bridge the gap between general-purpose language understanding and the stringent requirements of domain-specific tasks, enabling LLMs to perform effectively in highly specialized contexts.
Figure 1: Illustration of Growth Trends in Domain-Specific Knowledge Injection into LLMs. The chart displays the cumulative number of papers published between October 2022 and December 2024. Different colors and border styles represent various injection methods and domains, such as blue with a solid border denoting dynamic injection in the biomedical field.
Building on the foundational capabilities of general-purpose LLMs, knowledge injection techniques provide an effective means to address their limitations in handling specialized applications. Compared to the generalized approach of standard LLMs, knowledge injection offers two key advantages: 1) incorporating precise, domain-specific knowledge to improve accuracy and reliability in specialized tasks, and 2) allowing LLMs to dynamically adapt to new information or evolving knowledge bases, ensuring up-to-date expertise. These techniques bridge the gap between general-purpose understanding and domain-specific demands by leveraging both structured and unstructured knowledge sources. As a result, knowledge injection methods have been successfully applied in fields such as healthcare, chemistry, and legal analysis, significantly enhancing LLM performance. For example, biomedical LLMs  (Bolton et al., 2024; Yan et al., 2023) have demonstrated superior accuracy in tasks like medical diagnostics and regulatory compliance, while domain-specific models for material science (Xie et al., 2024; Antunes et al., 2024; Zhang et al., 2024a) have achieved advances in material property prediction and discovery. These dedicated models underscore the transformative potential of integrating domain knowledge into LLMs, unlocking solutions to complex, field-specific challenges.
Despite these advancements, early efforts in knowledge injection often treated domains independently, leading to a lack of standardization in methodologies and evaluation. As the volume of research continues to grow rapidly, with applications and studies proliferating across disciplines, the need for a comprehensive review becomes evident. This review aims to summarize the state of knowledge injection techniques, provide a systematic blueprint for future research, and identify key challenges, such as balancing scalability with domain-specific accuracy and enabling efficient, real-time knowledge updates.
To assist readers from various backgrounds in understanding knowledge injection techniques for LLMs and to complement existing surveys by addressing unresolved questions, we organize our survey paper as follows. After providing the foundational background knowledge in Section 2, we address a pivotal question: How can domain-specific knowledge be effectively integrated into LLMs? To answer this, we present a comprehensive framework for categorizing knowledge injection methods in Section 3. We delve into this topic by discussing: 1) Dynamic Knowledge Injection, which explains how external knowledge is retrieved and incorporated in real-time during inference to enhance reasoning capabilities; 2) Static Knowledge Embedding, which describes how domain knowledge is embedded into the model during training or fine-tuning to make it an inherent part of the model; 3) Adapters, which highlight modular techniques for storing and utilizing external knowledge without altering the primary model’s parameters; and 4) Prompt Optimization, which focuses on how carefully designed prompts enable models to utilize existing knowledge without requiring changes to their architecture. Another perspective for reviewing knowledge injection studies is their application across various domains. In Section 4, we categorize current application domains, such as materials science, chemistry, biology, and law. To guide the identification of appropriate tools and resources, we summarize commonly used benchmarks, open-source frameworks, and analyses on performance in Section 5. Based on these summaries, we discuss the challenges and opportunities in this evolving field in Section 6. The conclusions are summarized in Section 7.
2 Background
2.1 Domain-Specific Knowledge
Domain-specific knowledge refers to specialized information or expertise pertinent to a specific field or application, distinguishing it from general knowledge that spans across multiple domains. While general knowledge enables models to understand broad contexts, domain-specific knowledge is essential for addressing specialized tasks where precise, field-specific understanding is required. For instance, in scientific text processing (Bran et al., 2023), models must comprehend complex scientific terminologies, concepts, and methodologies to provide accurate and relevant answers. Similarly, in e-commerce search (Zhao et al., 2024a), understanding domain-specific terms such as product categories, technical specifications, or colloquial shopping language is crucial for delivering relevant search results and recommendations. In healthcare applications, LLMs must understand medical terminologies, diagnoses, treatment plans, and drug interactions. For example, biomedical question answering (Pei and others, 2024) and medical report summarization rely on integrating knowledge from medical literature like PubMed (Dernoncourt and Lee, 2017).
To address these needs, researchers have explored various methods for incorporating domain-specific knowledge into LLMs. In this paper, we aim to provide a survey of these various injection methods.
2.2 Knowledge Representation and Encoding
Knowledge can take different forms depending on the structure and application needs. For example, knowledge graphs (Zhang et al., 2024c) represent information as entities and relationships in a graph format, allowing structured reasoning and inference. These are widely used in tasks like question-answering and recommendation systems, where relationships between entities are crucial. Similarly, knowledge in text form, such as Wikipedia (Jeong et al., 2024), provides a vast corpus of unstructured information.
Knowledge can also be stored in vector space instead of readable text or graph formats. For instance, soft prompt tuning (Singhal et al., 2023a) learns useful knowledge in vector form, which is concatenated with the original input to guide LLMs in performing specific downstream tasks. In addition to external representations, knowledge can also emerge from within the model itself. For example, chain-of-thought prompting (Yao et al., 2024) introduces intermediate reasoning steps that help the model break down complex tasks into manageable parts. By explicitly reasoning through these steps, the LLM can utilize its internally stored information more effectively, resulting in better performance on tasks requiring logical reasoning, multi-step computation, or decision-making.
3 Paradigms of Knowledge Injection
This section identifies four major paradigms of knowledge injection: Dynamic Knowledge Injection, Static Knowledge Embedding, Adapters, and Prompt Optimization. These paradigms illustrate various mechanisms by which external domain-specific knowledge can be incorporated into LLMs. We utilize unified notations, as described in Table 1, to systematically represent the processes. External knowledge $\mathcal{K}$ is integrated into LLMs either by modifying the original parameters $\theta$, introducing additional parameters $\phi$, or leveraging auxiliary mechanisms.
Figure 2: Four knowledge injection paradigms for LLMs. (a) Dynamic Knowledge Injection retrieves external knowledge during inference for enhanced reasoning. (b) Static Knowledge Injection embeds external knowledge into model parameters during fine-tuning. (c) Modular Knowledge Adapters use plug-and-play modules to dynamically adapt to tasks or updates. (d) Prompt Optimization utilizes precise prompts to guide the LLM without altering its parameters.
3.1 Dynamic Knowledge Injection
Symbol Description
$\mathbf{x}$ Input to LLM
$\mathbf{y}$ Output of LLM
$M$ Backbone LLM Function
$\mathcal{K}$ External domain knowledge base
$\theta$ Parameters of LLM
$\phi$ Additional parameters introduced
$\mathcal{R}(\mathbf{x},\mathcal{K})$ Retrieval function fetches relevant elements of $\mathcal{K}$ given the input $\mathbf{x}$
$M(\mathbf{x};\theta)$ Represent LLM takes input $\mathbf{x}$ and produces an output, parameterized by $\theta$
$\Delta\theta$ Offsets to the original LLM’s parameters
Table 1: Summary of Symbols.
We define dynamic knowledge injection as the process of first retrieving information from external knowledge bases or knowledge graphs and then combining it with the input for use in LLMs:
$\mathbf{y}=M(\mathbf{x},\mathbf{\mathcal{R}(\mathbf{x},\mathcal{K})};\theta),$ (1)
where $\mathbf{x}$ represents the original input, $\mathcal{R}$ denotes the retrieval function, $\mathcal{K}$ is the external knowledge base, and $\theta$ are the model parameters, which remain unchanged. This paradigm offers several advantages, including ease of updating (hence the term ”dynamic injection”) and the ability to incorporate new knowledge without retraining the model. However, it also presents challenges, such as dependency on the quality of the knowledge base $\mathcal{K}$, the retrieval function $\mathcal{R}$, and limitations imposed by the maximum input length of the LLM.
3.2 Static Knowledge Embedding
Compared with dynamic knowledge retrieval, static knowledge embedding involves embedding knowledge into the model’s parameters through full or partial fine-tuning, making it less flexible to changes. Concretely, the model learns new parameters $\Delta\theta$ that encode domain knowledge from $\mathcal{K}$:
$\Delta\theta=\textstyle\arg\min_{\theta}\sum_{(\mathbf{x_{s}},\mathbf{y_{s}})\in\mathcal{K}}\mathcal{L}\bigl{(}M(\mathbf{x_{s}};\theta),\mathbf{y_{s}}\bigr{)},$
where $\mathcal{K}$ is the domain-specific knowledge base containing training samples $\mathbf{x_{s}}$ and $\mathbf{y_{s}}$, and $\mathcal{L}$ is a typical supervised training loss function. After optimization, the updated parameters $\Delta\theta$ are obtained.
At inference time, no further retrieval or external knowledge calls are required:
$\mathbf{y}=M\bigl{(}\mathbf{x};\Delta\theta\bigr{)}.$
This paradigm provides fast inference, as it eliminates additional retrieval steps and often results in stronger performance. However, it also comes with challenges, such as costly updates—requiring fine-tuning whenever domain knowledge changes—and scalability issues, as embedding large or frequently changing knowledge bases can be computationally expensive.
3.3 Modular Knowledge Adapters
To address the costly updates associated with static knowledge embedding, another paradigm, known as modular knowledge adapters, introduces small, trainable modules that can be inserted into or operate alongside the base model to store domain-specific knowledge while saving computational resources. In this approach, the original parameters $\theta$ of the LLM typically remain frozen, preserving the model’s general-purpose capabilities. Given a knowledge dataset $\mathcal{K}$, the adapter parameters $\phi$ are trained by minimizing the following objective:
$\phi=\textstyle\arg\min_{\phi}\sum_{(\mathbf{x_{s}},\mathbf{y_{s}})\in\mathcal{K}}\mathcal{L}\bigl{(}M(\mathbf{x_{s}};\theta,\phi),\mathbf{y_{s}}\bigr{)},$
where $M(\mathbf{x_{s}};\theta,\phi)$ represents the base model’s generation function enhanced with the new adapter parameters. At inference time, the enhanced model generates outputs as:
$\mathbf{y}=M\bigl{(}\mathbf{x};\theta,\phi\bigr{)}.$
This paradigm offers a parameter-efficient method to adapt LLMs to specific domains without modifying the original model weights. By freezing the base model’s parameters, the approach seeks to preserve previously acquired knowledge while enabling the seamless incorporation of new domain-specific information. However, this approach also introduces challenges, such as the need to design new architectural components and determine appropriate hyperparameters, including the size and number of adapters. These additional elements can increase the overall complexity of the model and its training process.
3.4 Prompt Optimization
Unlike previous approaches, prompt optimization does not retrieve knowledge from external sources. Instead, it focuses on fully leveraging or guiding the LLM to utilize its internal, pre-existing knowledge. The process can be formalized as:
$\mathbf{y}=M\bigl{(}[\mathbf{p},\mathbf{x}];\theta\bigr{)},$
where $\mathbf{p}$ represents a textual prompt containing implicit domain knowledge or specific instructions.
Prompt optimization offers significant advantages, including eliminating dependency on external domain knowledge bases and avoiding training. However, it also presents challenges, as designing effective prompts can be both complex and time-consuming. Additionally, long prompts may reduce the available context window, potentially affecting model efficiency and performance.
3.5 Comparison of the Four Paradigms
Paradigm Training Cost Inference Speed Limitations
Dynamic
Injection
None, but requires
retrieval module
Slower due to
retrieval latency
Relies heavily on
retrieval quality
Static
Embedding
High
(requires pretraining
or fine-tuning)
No extra cost
Fixed knowledge;
risks catastrophic
forgetting
Modular
Adapters
Low
(train small subset
of parameters)
Almost
unaffected
Sensitive to training
data quality
Prompt
Optimization
None
Almost
unaffected
Labor-intensive;
limited to pre-existing
knowledge
Table 2: Guidance on selecting knowledge injection paradigms based on training cost, inference speed, and limitations.
Dynamic knowledge injection integrates external knowledge at runtime, offering flexibility and adaptability to new information without increasing training cost. However, it requires an effective retrieval module, and the inference speed depends heavily on retrieval performance, which can slow down the overall process. Static knowledge embedding embeds domain expertise during pretraining or fine-tuning, requiring large-scale domain-specific data and significant training resources, including GPUs and time. While it incurs no additional inference cost, its limitations lie in the potential risks of catastrophic forgetting and its inability to adapt to evolving information. Modular adapters serve as a middle ground, allowing plug-and-play components to enhance domain-specific capabilities with minimal training data. Only a small subset of parameters needs to be trained, which reduces training costs, and inference speed is largely unaffected. However, the quality of training data significantly impacts the performance of this method. Prompt Optimization, on the other hand, avoids retraining entirely by leveraging carefully crafted inputs. It has no impact on inference speed but relies on extensive human effort to find optimal prompts. This method is limited in its ability to utilize new knowledge and primarily activates pre-existing knowledge.
We summarize these comparisons in Table 2 as a practical guide to help determine the most suitable method based on specific task requirements and scenarios.
4 Applications
4.1 Biomedicine
The biomedicine domain benefits from a wealth of specialized corpora, such as PubMed (Dernoncourt and Lee, 2017) and MedQA (Jin et al., 2021), enabling the development of LLMs specifically trained on biomedical texts. These models often follow the static knowledge embedding approach, leveraging the domain-specific richness of biomedical data. For instance, PMC-LLaMA (Wu et al., 2023) extends the LLaMA 7B model through further pretraining on 4.9 million PubMed Central articles curated from the S2ORC dataset (Lo et al., 2020), completing five epochs to embed biomedical knowledge effectively. Similarly, Med-PaLM 2 (Singhal et al., 2023b) builds on PaLM 2 via instruction fine-tuning. This fine-tuning incorporates a diverse mix of medical question-answering datasets, including MedQA, MedMCQA (Pal et al., 2022), and HealthSearchQA (Singhal et al., 2023a).
Beyond foundational models, integrating external tools and knowledge can further enhance performance. For example, GeneGPT (Jin et al., 2024) utilizes an LLM pretrained on code tasks to tackle GeneTuring tests by employing NCBI Web APIs. This approach combines in-context learning with an augmented decoding algorithm that identifies and executes API calls. Similarly, Med-PaLM (Singhal et al., 2023a) introduces vector prompts—representations that store and retrieve medical domain knowledge—to extend the capabilities of Flan-PaLM (Chung et al., 2024).
Domain Model Paradigms Knowledge Source Link
Biomedicine PMC-LLaMA (Wu et al., 2023) Static Knowledge Embedding PMC-OA, MedC-I, PubMedQA, MedMCQA, USMLE Link
Med-PaLM 2 (Singhal et al., 2023b) Static Knowledge Embedding MultiMed Link
DALK  (Li and others, 2024)
Dynamic Knowledge Injection
Prompt Optimization
MedQA, MedMCQA, MMLU, QA4MRE Link
ChronicCareGPT (Liu et al., 2024b) Prompt Optimization eRisk Link
SA-MDKIF (Xu et al., 2024b) Modular Knowledge Adapters MedQuA,emrQA, PubMedQA, MedQA \
MaLP (Zhang et al., 2024b) Modular Knowledge Adapters HealthCareMagic-100k, iCliniq Link
BioMedLM  (Bolton et al., 2024) Static Knowledge Embedding PubMed,MedMCQA,MedQA,MMLU,BioASQ Link
BiomedRAG (Li et al., 2024) Dynamic Knowledge Injection CHEMPROT,DDI,ade-corpus-v2,MTsample,ADInt,UMLS Link
MedINST (Han et al., 2024) Static Knowledge Embedding MedINST Link
Finance FLANG  (Shah and others, 2022) Static Knowledge Embedding
Financial PhraseBank,FiQA 2018 Task-1,
News Headline Classification, Named Entity Recognition,
Structure Boundary Detection,Question Answering
Link
BloomBergGPT (Wu and others, 2023)
Static Knowledge Embedding
Finance dataset (web, news, filings, press, Bloomberg),
Public dataset (the Pile, C4, Wikipedia)
FinMA  (Xie et al., 2023) Static Knowledge Embedding
FPB,FiQA-SA,Headline,NER,FinQA,
ConvFinQA,BigData22,ACL18,CIKM18
Link
FinGPT (Zhang et al., 2023) Modular Knowledge Adapters
Financial news, Company filings and announcements,
Social media discussions, Trends
Link
Fin-LLaMA  (Konstantinidis and others, 2024) Static Knowledge Embedding fin-llama-dataset Link
SNFinLLM (Zhao et al., 2024a)
Static Knowledge Embedding
FinEval, FinanceIQ,qEQA,FinC,KQA,MRC,cMRC
Materials ChemCrow (Bran et al., 2023) Dynamic Knowledge Injection 18 expert-designed tools Link
ChemDFM  (Zhao et al., 2024b) Static Knowledge Embedding
SciQ,PIQA,PubChem,ARC,USPTO
Link
ChemLLM (Zhang and others, 2024)
Static Knowledge Embedding
ChemData,ChemBench
Link
CrystaLLM  (Antunes et al., 2024)
Static Knowledge Embedding
Materials Project, OQMD, NOMAD
Link
ScholarChemQA  (Chen and others, 2024)
Static Knowledge Embedding
AG News,Yahoo Answers ,Yelp-5,Amazon-5 Link
DARWIN 1.5 (Xie et al., 2024) Static Knowledge Embedding FAIR datasets Link
ChemAgent (Tang and others, 2025)
Dynamic Knowledge Injection
Prompt Optimization
Google API, Docker container, Internet,
Hardware API document, Physical world hardware
Link
Mental Health MeChat (Qiu et al., 2023) Dynamic Knowledge Injection SMILECHAT, PsyQA Link
MindChat (Xin Yan, 2023) Static Knowledge Embedding Multi-turn psychological dialogue data Link
SoulChat (Chen et al., 2023) Static Knowledge Embedding Long-text counseling sessions Link
EmoLLM (Yang et al., 2024)
Static Knowledge Embedding
Modular Knowledge Adapters
CPsyCounD Link
Education EduChat (Dan et al., 2023) Static Knowledge Embedding
Textbooks Data, Open QA Data,
Emotional Support Data, Socratic Teaching Data
Link
QiaoBan (Weixiang et al., 2023) Prompt Optimization Children’s emotional education dialogue data Link
HiTA (Liu et al., 2024a) Dynamic Knowledge Injection Educator curated database \
SocraticLM (Liu et al., 2024c) Modular Knowledge Adapters SocraTeach dataset \
CyberQ (Agrawal et al., 2024)
Static Knowledge Embedding
Dynamic Knowledge Injection
AISecKG, Q&A \
Social Science SocialLLM (Jiang and Ferrara, 2023)
Static Knowledge Embedding
Prompt Optimization
Covid-Political, Election2020, COVID-Morality,
Ukr-Rus-Suspended, Ukr-Rus-Hate,
Immigration-Hate-08, Immigration-Hate-05
FPS (Liu et al., 2024e) Prompt Optimization Fake News Dataset, Big Five Personality Traits Link
FUSE (Liu et al., 2024f) Prompt Optimization True News Dataset, Big Five Personality Traits \
Table 3: Summary of the domain-specific knowledge injection studies. We categorize current work according to their research domain and knowledge injection method.
4.2 Finance
Fine-tuned financial LLMs have demonstrated significant advancements in adapting general-purpose models to domain-specific tasks through task-specific training. PIXIU (Xie et al., 2023) fine-tunes LLaMA on 136K instruction samples tailored to financial tasks, enabling the model to handle a wide range of domain-relevant scenarios. Instruct-FinGPT (Zhang et al., 2023) fine-tunes LLaMA on 10K instruction samples derived from two financial sentiment analysis datasets, focusing primarily on finance classification tasks. FinGPT (Yang et al., 2023) introduces a comprehensive end-to-end framework for training and deploying FinLLMs in the financial industry. Utilizing the LoRA technique, FinGPT fine-tunes open-source LLMs like LLaMA and ChatGLM with approximately 50K task-specific samples, achieving efficient fine-tuning without full model retraining.
In contrast, scratch-trained financial LLMs aim to create models specifically designed for financial tasks from the ground up. BloombergGPT (Wu and others, 2023) leverages a subset of 5 billion tokens from Bloomberg-specific data, representing only 0.7% of its total training corpus, to tailor its model to financial applications. XuanYuan 2.0 (Zhang and Yang, 2023) combines 366 billion tokens for pre-training with an additional 13 billion tokens for fine-tuning, creating the largest Chinese financial chat model. Similarly, Fin-T5 (Lu et al., 2023) introduces a Chinese financial pre-training language model built on the T5 architecture, utilizing a 300GB financial corpus. Furthermore, SNFinLLM (Zhao et al., 2024a) dynamically incorporates real-time financial data during inference to enhance decision-making capabilities, demonstrating the value of domain-specific pretraining and adaptability in financial LLMs.
4.3 Materials
In contrast to biomedicine, where significant efforts have been devoted to static knowledge embedding due to the availability of extensive corpora, research in materials and chemistry has primarily focused on utilizing task-related tools that align with the dynamic knowledge injection paradigm.
For instance,  Xie et al. (2024) demonstrated how Darwin 1.5 leverages natural language inputs and a two-stage training strategy to achieve significant improvements in materials discovery and design tasks. Bran et al. (2023) introduced ChemCrow, a framework that augments LLMs with chemistry-expert-designed tools for downstream tasks such as organic synthesis and drug discovery. There are also studies on prompt optimization Tang and others (2025), demonstrating that better-designed planning prompts can effectively utilize the model’s internal knowledge to orchestrate complex tasks. This approach capitalizes on the planning and execution capabilities of multiple LLMs to achieve autonomy in chemical experimentation.
More recently, there has been growing interest in exploring static knowledge embedding and modular knowledge adapters within the chemistry domain. For example, Chen and others (2024) curated a QA dataset to fine-tune pretrained models like BERT and LLMs such as Llama, aiming to enhance their performance in chemistry-related tasks. Similarly, Xie et al. (2024) introduced Darwin 1.5, an open-source large language model tailored for materials science.
4.4 Human-Centered Science
The last domain we introduce is human-centered science, which encompasses a wide range of applications such as psychological counseling, financial forecasting, social behavior prediction, and legal reasoning. All these domains center on understanding and interacting with human needs, behaviors, and decision-making processes.
In mental health, datasets like PsyQA (Sun et al., 2021) provide a foundation for training models in psychological counseling scenarios. For instance, SoulChat (Chen et al., 2023), a model fine-tuned on over 100,000 long-text counseling sessions using static knowledge embedding, is designed for empathic conversations. Similarly, MeChat (Qiu et al., 2023) employs dynamic knowledge injection to adapt to real-time inputs, significantly enhancing its emotional support capabilities. These advancements demonstrate the potential of human-centered science in addressing complex, real-world challenges through personalized and context-aware solutions.
In the education domain, LLMs have shown immense potential in addressing challenges such as personalized learning, curriculum alignment, and interactive teaching. Personalized learning, for example, requires models to adapt to individual needs, providing tailored feedback and emotional support. EduChat (Dan et al., 2023) tackles this by leveraging educational theories from psychology and pedagogy through static knowledge embedding, enabling tasks like open Q&A, composition correction, and emotional support. Similarly, QiaoBan (Weixiang et al., 2023) focuses on child-centered education by using prompt optimization to adapt the model’s behavior based on child psychology and emotional well-being, catering specifically to young learners. Domain-specific education and interactive teaching have also seen advancements through LLMs. CyberQ (Agrawal et al., 2024) blends static knowledge embedding and dynamic knowledge injection via AISecKG (Agrawal, 2023), generating Q&A based on cybersecurity best practices. Interactive teaching, on the other hand, benefits from models like SocraticLM (Liu et al., 2024c), which employs adapters fine-tuned on the SocraTeach dataset to engage students in critical thinking and problem-solving.
For social sciences, models like SocialLLM (Jiang and Ferrara, 2023) combine static knowledge embedding and dynamic knowledge injection to analyze human behavior in social networks. Adapters facilitate large-scale data integration while prompt optimization guides the model to focus on specific social behavior patterns. Models like FPS (Liu et al., 2024e) and FUSE (Liu et al., 2024f) use prompt optimization to simulate the spread and evolution of fake news in social networks, helping understand misinformation’s impact.
A summary of the mainstream models and their information is provided in Table 3. More models across various domains can be found at: Survey-official-repo.
5 Tools, Resources, and Analysis
5.1 Knowledge Injection Framework
In this section, we provide a detailed introduction to four open-source frameworks categorized under different knowledge injection methods to facilitate understanding and application: KnowGPT (Zhang et al., 2024c) for Dynamic Knowledge Injection, StructTuning (Liu et al., 2024d) for Static Knowledge Embedding, K-Adapter (Wang et al., 2021) for Modular Knowledge Adapters, and SelfLift (Cheng et al., 2024) for Prompt Optimization.
KnowGPT dynamically combines knowledge graphs with prompt optimization by leveraging reinforcement learning to extract highly relevant subgraphs from the knowledge graph. These subgraphs are represented as triples and transformed into natural language prompts that language models can interpret and utilize via diverse prompt templates. The KnowGPT framework significantly reduces the API call costs of LLMs while enhancing their performance in domain-specific tasks.
StructTuning uses a structure-aware approach to embed domain knowledge into pre-trained models with a two-stage strategy: Structure-Aware Continual Pre-Training encodes knowledge into the model’s parameters, and Structure-Aware Supervised Fine-Tuning refines understanding through structured QA tasks. This framework demonstrates significant performance improvements in knowledge-driven tasks such as relation classification and question answering, achieving a balance between generality and efficiency.
K-Adapter stores knowledge within adapter modules. Its core method involves freezing the original model parameters and assigning an independent, task-specific adapter for each type of knowledge. These adapters are inserted as independent modules into the intermediate layers of the model to generate enhanced representations of specific knowledge. This design effectively mitigates the issue of catastrophic forgetting, preventing newly injected knowledge from overwriting the model’s pre-existing knowledge.
Finally, SelfLift iteratively employs a retrieval-augmented generator to create an unbounded memory pool and uses a memory selector to choose one output as memory for the subsequent generation round. This is an excellent demonstration of prompt optimization, where the model’s outputs are dynamically refined and reused to enhance its overall performance and coherence in subsequent tasks.
5.2 Datasets and Benchmarks
We summarize commonly used datasets or benchmarks for domain-specific LLM study in Table 3, observing significant variation in dataset richness across domains. Biomedicine boasts numerous high-quality datasets, such as PubMed, PubMedQA (Jin et al., 2019), and BioASQ Tsatsaronis and others (2012), that support tasks such as question answering and clinical summarization. In contrast, materials and chemistry have more limited resources, and datasets like USPTO and Enzymes focus on chemical reactions. Miscellaneous datasets are scattered across other domains, such as PsyQA and SmileChat in mental health, SocraTeach, and Children’s emotional education dialogue data dataset in education. This diversity underscores the effort to tailor LLMs to specialized fields while emphasizing the need for broader curation of benchmarks in underrepresented domains.
5.3 Performance Comparison of Domain-specific LLM and General-domain LLM
Model Domain Type Size MedQA PubMedQA MedMCQA
Med-Gemini Specific 91.1 - -
GPT-4 General 90.2 80.4 73.7
Med-PaLM 2 Specific 85.4 81.8 72.3
PMC-LLaMA Specific 13B 56.3 77.9 56.0
BioMedLM Specific 2.7B 50.3 74.4
Llama 2 General 70B 43.7 74.3 35.0
Galactica General 120B 44.4 77.6 52.9
Table 4: Performance comparison of domain-specific and general-domain model performance on medical benchmarks.
Since there are also general-domain strong LLMs, it is necessary to discuss the comparison between domain-specific LLMs and general ones to determine if a specific knowledge injection process is essential. Here, we take the biomedical domain as an example due to significant research efforts in this field, as shown in Table 4. The results are collected from corresponding papers or paperswithcode.com.
First, we can observe that closed-source LLMs are currently the most effective models, while the performance gap between general-domain and domain-specific LLMs is relatively narrow. For example, both GPT-4 and Med-Gemini (Saab et al., 2024) achieve outstanding performance, with scores higher than 90 on the MedQA dataset. However, due to the lack of transparency in closed-source LLMs, open-source LLM efforts should not be overlooked. In this field, domain-specific LLMs often outperform general-domain models. For instance, PMC LLaMA-13B outperforms LLaMA2-70B by more than 10 points on the MedQA dataset. This demonstrates the value of domain-specific LLMs in achieving superior performance on specialized tasks. While general-domain models can deliver strong results, incorporating domain-specific knowledge allows for significant improvements, particularly in open-source initiatives. This underscores the importance of investing in domain-specific LLMs to address the unique challenges of specialized fields.
6 Challenges and Opportunities
6.1 Integrated Knowledge Consistency
Knowledge injection allows LLMs to incorporate and integrate different domain-specific knowledge. However, retrieved knowledge may conflict with the model’s pre-trained representations or other retrieved facts, leading to inconsistencies in outputs (Xu et al., 2024a). For example, in healthcare or legal analysis, conflicting treatment protocols or contradictory legal precedents could arise, resulting in unreliable decisions and undermining the system’s trustworthiness.
To address this, future research must focus on detecting inconsistencies, resolving conflicts, and maintaining consistency in integrated knowledge. Conflicts can be addressed by prioritizing reliable sources, applying domain-specific rules, or using ensemble techniques to balance multiple perspectives. Alignment algorithms and validation modules can further ensure that retrieved knowledge aligns with the model’s reasoning processes and is reliable before influencing outputs. These efforts are essential to enhance the reliability and applicability of knowledge-enhanced LLMs in complex, high-stakes domains.
6.2 Cross-Domain Knowledge Transfer
Cross-domain knowledge transfer involves equipping LLMs with the ability to generalize knowledge across diverse and distinct fields. While this significantly expands their applicability, it also introduces challenges due to the complexity and diversity of domain-specific terminologies, ontologies, and reasoning patterns. For example, transferring knowledge from chemistry to healthcare might require reconciling differing data structures and reasoning frameworks.
Overcoming these challenges necessitates advancements in modular knowledge representation and transfer learning techniques. Future efforts could explore hybrid approaches that blend static embeddings with dynamic retrieval, enabling LLMs to adapt knowledge flexibly across domains without compromising depth. Additionally, creating standardized cross-domain benchmarks and datasets could facilitate systematic evaluation and foster innovation in multi-domain knowledge transfer methodologies.
7 Conclusion
LLMs enhanced by domain-specific knowledge have shown remarkable potential and garnered increasing research interest. This survey systematically reviews LLM knowledge injection systems, exploring knowledge representation methods, integration strategies, and mechanisms for preserving model generality. We also summarize applications across biomedicine, chemistry, and computational social science domains. By highlighting standard datasets, benchmarks, challenges, and future opportunities, we aim to provide a valuable resource that inspires further exploration of knowledge-enhanced LLMs for domain-specific challenges.